Quantitative Problem 1: Suppose that 1 Swedish krona could be purchased in the foreign exchange market today for $0.4. If the krona appreciated 9% tomorrow against the dollar, how many kronas would a dollar buy tomorrow? Round your answer to 2 decimal places. Do not round intermediate calculations.
kronas

Quantitative Problem 2: Suppose the exchange rate between the U.S. dollar and the South African rand was 10 rand = $1 and the exchange rate between the U.S. dollar and the Israeli shekel was 1 shekel = $0.35. What was the exchange rate between the South African rand and the Israeli shekel? Round your answer to 2 decimal places. Do not round intermediate calculations.
rands per shekel
Answer #1

1.

1 krona = $0.4.

After appreciation, 1 krona = $0.4 (1+0.09) = $0.436.

So, $1 = (1/0.436) krona

$1 = 2.29.

1 dollar will buy 2.29 Swedish kronas.

2.

$1 = 10 rand.

1 shekel = $0.35.

Divide the above equation by 0.35.

2.86 shekel = $1.

10 rands = 2.86 shekel.

Exchange rate between rand and shekel (rand/shekel) = (10/2.86) = 3.50.